OpenStudy (anonymous):

simplify sqrt 3c^2d (multiplied by) sqrt 6cd^2

OpenStudy (watchmath):

\(\sqrt{3c^2d\cdot 6cd^2}=\sqrt{2\cdot 9\cdot c^2\cdot d^2\cdot c}=3cd\sqrt{2c} \)